OpenStudy (anonymous):

Your roommate is working on his bicycle and has the bike upside down. He spins the 50.0 -diameter wheel, and you notice that a pebble stuck in the tread goes by three times every second. What is the speed and acceleration of the pebble?

OpenStudy (anonymous):

the pebble has 3 rps or 180 rpm(rotation per minute) and its angular speed will be 2*pi*rps=18.84 rad/s..hence the speed will be angular speed*radius=18.84*25/100(if diameter is in centimeters) =4.71 m/s & acceleration = speed^2/radius=88.73 m/s^2